OpenStudy (anonymous):

use rstional exponents to simplify 4sqrtx^8

OpenStudy (anonymous):

\[\sqrt[4]{x ^{8}?}\]

OpenStudy (hunus):

\[\sqrt[a]{x^b}\] is the same as \[x^{b/a}\] So given \[\sqrt[4]{x^8}\] It is the same as \[x^2\]

OpenStudy (anonymous):

so my answer would be x^2

OpenStudy (hunus):

Yes, sorry for the late reply